An ideal gas at temperature 50 degrees C is confined to a box of volume 0.1 m^3.

a. What is the average kinetic energy of the particles in the gas?

b. If there are 10^25 particles in the gas, how much energy does it contain?

c. If there are 10^25 particles in the box, what is the pressure of the gas?
Answer #1

la) MIN (6) KE. - 3x 8-14 x (50+273) 26.02 x 1023 = 6.69 x 10-21 J Answer E= 1025 x 6.69€ 10-21 = 6.69 x 104 J Ansamer PrenRI